OpenStudy (anonymous):

Determine of (1/2,-4) is a solution to 4x-3y=-10.Please show all of your work.

OpenStudy (unklerhaukus):

does\[(x,y)=(1/2,-4)\]satisfy\[4x-3y=-10\]? \[4(1/2)-3(-4)\]\[=2+12\]\[=14\]\[≠-10\] the point does not satisfy the equation, ie the point is not on the line